A researcher compares the effectiveness of two different instructional methods for teaching anatomy. A sample of 59 students using Method 1 produces a testing average of 85.5. A sample of 31 students using Method 2 produces a testing average of 74.4. Assume that the population standard deviation for Method 1 is 19, while the population standard deviation for Method 2 is 7.55. Determine the 98% confidence interval for the true difference between testing averages for students using Method 1 and students using Method 2.

Step 2 of 3:

Calculate the margin of error of a confidence interval for the difference between the two population means. Round your answer to six decimal places.

Step 3 of 3:

Construct the 98% confidence interval. Round your answers to one decimal place.
Answer #1


Pooled Variance
sp = sqrt(s1^2/n1 + s2^2/n2)
sp = sqrt(361/59 + 57.0025/31)
sp = 2.820892

Given CI level is 0.98, hence α = 1 - 0.98 = 0.02
α/2 = 0.02/2 = 0.01, tc = t(α/2, df) = 2.326

Margin of Error
ME = tc * sp
ME = 2.326 * 2.820892
ME = 6.561395

CI = (x1bar - x2bar - tc * sp , x1bar - x2bar + tc * sp)
CI = (85.5 - 74.4 - 2.326 * 2.820892 , 85.5 - 74.4 - 2.326 * 2.820892
CI = (4.5 , 17.7)